# On Limiting Probability Distributions of Higher Order Markov Chains

### v1

@@ Abstract
The limiting (or stationary) probability distribution is one of the key characteristics of a Markov chain since it shows its long-term behavior. In this paper, for a higher order Markov chain, we establish a sufficient condition for the existence of its limiting probability distribution. This condition is built upon the regularity of its transition tensor. Our results extend the corresponding conclusions for first order chains. Besides, they complement the existing results concerning higher order chains that rely on approximation schemes or two-phase power iterations.
@@ 1 Introduction
Let $X=\{X_{t}:t=1,2,\ldots\}$ be a stochastic process, where $X_{t}$ is a random variable representing the state of a system at time $t$. Assume that $X$ takes values in the state space $S=\{1,2,\ldots,n\}$. Then, $X$ is called a first order Markov chain if the following Markovian property holds:
$\Pr(X_{t+1}=i|X_{t}=j,X_{t-1}=k_{1},\ldots,X_{1}=k_{t-1})=\Pr(X_{t+1}=i|X_{t}=j)$
for any $t\geq 1$ and $i,j,k_{1},\ldots,k_{t-1}\in S$. This property can be interpreted as that the future evolution of the process $X_{t+1}$ depends on its current status $X_{t}$ only.
We also assume that the above chain $X$ is homogeneous, meaning that $\Pr(X_{t+1}=i|X_{t}=j)$ is independent of $t$. Under this assumption, we denote
$\Pr(X_{t+1}=i|X_{t}=j)=p_{ij},$
where $i,j\in S$. These are the transition probabilities of the chain. Accordingly, $P=[p_{ij}]\in\mathbb{R}^{n\times n}$ is the transition matrix, which is column stochastic, namely $0\leq p_{ij}\leq 1$ for any $i,j\in S$ and $\sum_{i=1}^{n}p_{ij}=1$ for any $j\in S$.
The preceding Markovian property can be adapted in a way so that the future evolution of the process depends not only on its current status but also on one or more of its immediately past statuses. Formally, $X$ is called an $(m-1)$th order, where $m\geq 2$, Markov chain if the Markovian property takes the following $(m-1)$th order form:
@@
for any $t\geq m-1$ and $i_{1},\ldots,i_{m},\ldots,i_{t+1}\in S$.
Again, under the homogeneity assumption, we denote
@@
where $t\geq m-1$ and $i_{1},i_{2},\ldots,i_{m}\in S$, and call it the transition probability from states $(i_{2},\ldots,i_{m})$ to state $i_{1}$. Moreover, the resulting $m$th order, $n$-dimensional tensor ${\cal P}=[p_{i_{1}i_{2}\ldots i_{m}}]$ is the transition tensor, which is “column” stochastic, namely $0\leq p_{i_{1}i_{2}\ldots i_{m}}\leq 1$ for any $i_{1},i_{2},\ldots,i_{m}\in S$ and $\sum_{i_{1}=1}^{n}p_{i_{1}i_{2}\ldots i_{m}}=1$ for any $i_{2},\ldots,i_{m}\in S$. By the term higher order chains, we refer to the case when $m\geq 3$. Clearly, when $m=2$, the chain reduces to a first order one.
For the past decade or so, higher order Markov chains have drawn much attention in the literature, see [1, 2, 3, 4, 7, 8, 9, 10, 11, 12, 13, 20, 21, 26]. A focal point in a majority of these existing works is the limiting probability distribution of a higher order chain. The following definitions, however, includes the first order case as well since it is also relevant here.
@@
Assume $m\geq 2$. For an $(m-1)$th order (homogeneous) Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S=\{1,2,\ldots,n\}$,
@@
Under the assumptions of Definition 1, if $\lim_{t\rightarrow\infty}x_{t}$ exists and, moreover, it does not depend on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$, then
@@
We observe that $\pi$ itself is a probability distribution, namely $\pi\geq 0$ and $\sum_{i=1}^{n}\pi_{i}=1$. It is an important characteristic of the chain since it provides useful information regarding its long-term behavior. It should be mentioned that $\lim_{t\rightarrow\infty}x_{t}$ may not exist or may depend on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$. An example about this shall be provided later after necessary background material is introduced.
For a first order chain with transition matrix $P$, a well-known sufficient condition regarding the existence of the limiting probability distribution is $P$ being regular. Denote $P^{k}=[p^{(k)}_{ij}]\in\mathbb{R}^{n\times n}$, where $k=1,2,\ldots$. Then, $p^{(k)}_{ij}$ represents the $k$-step transition probability from states $j$ to $i$, that is,
@@
for any $t\geq 1$. The transition matrix $P$ is called regular if there exists some $k\geq 1$ such that $P^{k}>0$. Under such a condition, $\lim_{t\rightarrow\infty}x_{t}=\pi$ exists and is independent of the initial probability distribution of $X_{1}$, and, in addition, $\pi>0$. Moreover, $\lim_{k\rightarrow\infty}P^{k}=\pi e^{T}$, where $e\in\mathbb{R}^{n}$ is the vector of all ones. For these existing results concerning regular first order chains, we refer the reader to [14, 17].
It cannot be overemphasized that regardless of the order of the chain, its limiting probability distribution, if it exists, does not depend on where the chain starts. In the first order case, this limiting probability distribution is also called the stationary probability distribution. The latter term has been used interchangeably in the literature on higher order chains. We, however, choose the former since for a higher order chain, this distribution lacks the interpretation of stationarity when it comes to the evolution of the chain.
In light of the aforementioned results on first order chains, it is natural to raise the question of whether similar results can be established on higher order chains. As far as we can see, this question has thus far been partially answered by certain approximations of the limiting probability distribution, a two-phase procedure under a stronger regularity requirement, or a matrix framework that only deals with second order chains. In this paper, we shall approach the question here largely from a tensor formulation perspective to provide a more complete answer.
Before proceeding, it should be pointed out that due to consideration of computational tractability, a vast majority of existing works on the limiting probability distribution of a higher order chain have taken the so-called $Z$-eigenvector approach [21] — see a brief survey in Section 3. This approach, however, provides only an approximation of the limiting probability distribution [5, 26]. Moreover, such an approximation, in general, does not reflect the true nature of the limiting probability distribution [26]. In comparison, our results here concern the exact limiting probability distribution.
2 Tensor Formulation and Matricization
The concept of higher order Markov chains dates back to as early as 1930s [16]. Most, if not all, classical references [14, 15], however, only talk briefly about their connection with first order chains but do not provide detailed discussion. In the recent works we have cited in the introductory section, a tensor framework has emerged as the main instrument for studying higher order chains. Nevertheless, results regarding such a framework appear to be a bit scattered, particularly relevant to the limiting probability distribution problem. In this section, therefore, we first set the stage methodically.
Assuming $m\geq 3$, we continue with an $(m-1)$th order (homogeneous) Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S=\{1,2,\ldots,n\}$, with an $m$th order, $n$-dimensional transition tensor ${\cal P}=[p_{i_{1}i_{2}\ldots i_{m}}]$.
Similar to the first order case, the $k$-step, with $k\geq 1$, transition probability from states $(i_{2},\ldots,i_{m})$ to state $i_{1}$ is given by
@@ 2 Tensor Formulation and Matricization
These $k$-step transition probabilities form a stochastic tensor too, which is denoted by ${\cal P}^{(k)}=[p^{(k)}_{i_{1}\ldots i_{m}}]$ and has the same size as $\cal P$. The special case of $k=0$ leads to the so-called identity tensor ${\cal I}=[\delta_{i_{1}i_{2}\ldots i_{m}}]$ — note that this identity tensor is different from those in [18, 23]. In addition, it is known that the $k$-step transition probabilities can be determined by the following recurrence formula [1, 9]:
$p^{(k+1)}_{i_{1}i_{2}\ldots i_{m}}=\sum_{j=1}^{n}p_{ji_{2}\ldots i_{m}}p^{(k)}_{i_{1}ji_{2}\ldots i_{m-1}},~{}k=0,1,2,\ldots.$ (2.1)
In [10], (2.1) is written as a tensor product ${\cal P}^{(k+1)}={\cal P}\boxtimes{\cal P}^{(k)}$. In this sense, ${\cal P}^{(k)}$ can be thought of as the $k$th power of $\cal P$. The identity tensor, on the other hand, satisfies ${\cal P}\boxtimes{\cal I}={\cal P}$. The operation $\boxtimes$ is distributive, yet it is neither commutative nor associative.
Next, we recall the concepts of irreducibility, ergodicity, and regularity when it comes to the higher order chain $X$. These can be extended to nonnegative tensors too, but we shall restrict ourself to the stochastic scenario here. Specifically, the transition tensor $\cal P$ is said to be :
@@
Obviously, regularity implies ergodicity. In the meantime, it is also known that ergodicity implies irreducibility but the reverse does not hold [9]. The latter is one of the differences between the higher order and the first order cases since irreducibility and ergodicity are the same on a first order chain.
We now review a systematical way the higher order chain $X$ can be reduced to a first order one. Denote $T=\{i_{1}i_{2}\ldots i_{m-1}:i_{1},i_{2},\ldots,i_{m-1}\in S\}$. It is the collection of multi-indices of length $m-1$. These multi-indices are arranged via linear indexing [22]. When $m=4$ and $n=2$, for example,
@@
Clearly, the cardinality of $T$ is $N=n^{m-1}$. Consider random variables $Y_{t}=(X_{t},X_{t-1},\ldots,X_{t-m+2})$ that take values in $T$. When
$X_{t}=i_{1},X_{t-1}=i_{2},\ldots,X_{t-m+2}=i_{m-1},$
for example, we have $Y_{t}=i_{1}i_{2}\ldots i_{m-1}$. Let $Y=\{Y_{t}:t=m-1,m,\ldots\}$ be the first order chain on state space $T$. It is also referred to as the reduced first order chain obtained from $X$. The transition matrix $Q\in\mathbb{R}^{N\times N}$ of $Y$ is given by
@@
We observe that $q_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=p_{i_{1}i_{2}\ldots i_{m-1}j_{m}}$ when $i_{\ell}=j_{\ell}$ for all $\ell=2,3,\ldots,m-1$; otherwise, $q_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=0$. Again, considering $m=4$ and $n=2$, we arrive at
@@
In general, $p_{i_{1}i_{2}\ldots i_{m}}$ is located in $Q$ on row
$i_{1}+n(i_{2}-1)+\ldots+n^{m-2}(i_{m-1}-1)$
and column
$i_{2}+n(i_{3}-1)+\ldots+n^{m-2}(i_{m}-1).$
We shall give an easier way of constructing $Q$ later in this section.
As in (1.3), the $k$-step transition probability of $Y$ from states $j_{2}j_{3}\ldots j_{m}$ to $i_{1}i_{2}\ldots i_{m-1}$ is given by
@@
It should be pointed out that problems regarding the higher order chain $X$ cannot always be fully addressed by resorting to the reduced first order chain $Y$. To explain this, we mention several examples here: (1) $Y$ may be non-ergodic even though $X$ is ergodic [10], (2) the mean first passage times of $X$ may not be obtained in a direct fashion from those of $Y$ [10], and (3) a recurrent state in $X$ may no longer be a recurrent state in $Y$ [15]. The reduced first order chain, nevertheless, can still be a useful tool as can be seen shortly.
Moving on, let $P\in\mathbb{R}^{n\times N}$ be the mode-$1$ matricization [19] of the transition tensor $\cal P$. Simply put, $P$ consists of the frontal slices of $\cal P$ which are arranged side by side and ordered according to linear indexing. Taking the case when $m=4$ and $n=2$, $P$ has the form
@@
We can refer to the entries of $P$ using multi-indices in the form of $p_{i_{1},i_{2}\ldots i_{m}}$, which is the same as $p_{i_{1}i_{2}\ldots i_{m}}$. In general, $p_{i_{1}i_{2}\ldots i_{m}}$ is located in $P$ on row $i_{1}$ and column
$i_{2}+n(i_{3}-1)+\ldots+n^{m-2}(i_{m}-1).$
Similarly, $P^{(k)}$ stands for the mode-$1$ matricization of ${\cal P}^{(k)}$. By convention, $P^{(0)}$ is the mode-$1$ matricization of the identity tensor $\cal I$, that is,
@@
Incidentally, the transition matrix $Q$ of the first order chain $Y$ is constructed in [26] in a quite complicated way by using the product of a block diagonal matrix and a permutation matrix. By the mode-$1$ matricization of ${\cal P}^{(k)}$, the construction of $Q$ can be simplified as
@@
The above mode-$1$ matricization of $\cal P$ can bridge the probability distribution of the higher order chain $X$ to that of the reduced first order chain $Y$. Specifically, let $y_{t}=x_{t,t-1,\ldots,t-m+2}\in\mathbb{R}^{N}$ be the probability distribution of $Y$ at time $t$, whose $i_{1}i_{2}\ldots i_{m-1}$th entry is given by
@@
for any $i_{1}i_{2}\ldots i_{m-1}\in T$. As usual, these entries are ordered again by linear indexing. As an example, with $m=4$ and $n=2$, we obtain
@@
By $x_{t}$ in (1.2) and $y_{t}$ as above, it can be shown [21] that
@@
Formula (2.5) is exactly what the two-phase procedure in [26] is based on. Besides, both $x_{t}$ and $y_{t}$ are further formulated as tensors in [26].
With the background material we have presented thus far, it is time to give an example to illustrate that, in general, $\lim_{t\rightarrow\infty}x_{t}$ may depend on the initial probability distribution of $X_{1},X_{2},\ldots,X_{m-1}$.

### v2

@@ Abstract
The limiting probability distribution is one of the key characteristics of a Markov chain since it shows its long-term behavior. In this paper, for a higher order Markov chain, we establish some properties related to its exact limiting probability distribution, including a sufficient condition for the existence of such a distribution. Our results extend the corresponding conclusions on first order chains. Besides, they complement the existing results concerning higher order chains which rely on approximation schemes or two-phase power iterations. Several illustrative example are also given.
@@ 1 Introduction
Let $m\geq 2$ be an integer. An $(m-1)$th order Markov chain is a stochastic process $X=\{X_{t}:t=1,2,\ldots\}$, where each $X_{t}$ takes values in state space $S=\{1,2,\ldots,n\}$, such that
@@
for all $t\geq m-1$ and $i_{1},\ldots,i_{m},\ldots,i_{t+1}\in S$. Throughout this paper, the chain $X$ is assumed to be homogeneous, meaning that the probability as in (1.1) is independent of $t$. Hence, we can denote
@@
for any $t\geq m-1$ and $i_{1},i_{2},\ldots,i_{m}\in S$. This is the transition probability from states $(i_{2},\ldots,i_{m})$ to state $i_{1}$. The resulting $m$th order, $n$ dimensional tensor ${\cal P}=[p_{i_{1}i_{2}\ldots i_{m}}]$ is called the transition tensor. Since $0\leq p_{i_{1}i_{2}\ldots i_{m}}\leq 1$ for any $i_{1},i_{2},\ldots,i_{m}\in S$ and $\sum_{i_{1}=1}^{n}p_{i_{1}i_{2}\ldots i_{m}}=1$ for any $i_{2},\ldots,i_{m}\in S$, $\cal P$ is called stochastic. The chain is said to be higher order whenever $m\geq 3$. In the first order case, the transition tensor $\cal P$ is often denoted by $P\in\mathbb{R}^{n\times n}$, which is called the transition matrix .
For the past decade or so, higher order Markov chains have drawn much attention in the literature, see [1, 2, 3, 4, 7, 8, 9, 10, 11, 13, 14, 20, 21, 26]. A focal point in a majority of these existing works is the limiting probability distribution of a higher order chain. The following definitions, however, include the first order case too since it is also relevant here.
@@
For an $(m-1)$th order, where $m\geq 2$, Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S=\{1,2,\ldots,n\}$,
@@
Under the assumptions of Definition 1, if $\lim_{t\rightarrow\infty}x_{t}$ exists and if this limit does not depend on the initial probability distributions, i.e., $x_{1},x_{2},\ldots,x_{m-1}$, then
@@
Clearly, $\pi$ satisfies $\pi\geq 0$ and $\sum_{i=1}^{n}\pi_{i}=1$ and thus is itself a probability distribution. It is an important characteristic of the chain since it provides useful information regarding its long-term behavior. The limit $\lim_{t\rightarrow\infty}x_{t}$, however, may not exist and may depend on the initial probability distributions. An example for this shall be given later after we introduce some necessary background material. It is natural, therefore, to ask under what conditions the limiting probability distribution exists and what properties it must satisfy.
For a first order chain with transition matrix $P$, a well-known sufficient condition regarding the existence of the limiting probability distribution is $P$ being regular. Denote [^1] We may also denote this as $P^{(k)}$. $P^{k}=[p^{(k)}_{ij}]\in\mathbb{R}^{n\times n}$, where $k=1,2,\ldots$. This $p^{(k)}_{ij}$ represents the $k$-step transition probability from states $j$ to $i$, that is,
@@
for all $t\geq 1$. Then, the transition matrix $P$ (as well as the chain) is called regular if there exists some $k\geq 1$ such that $P^{k}>0$. Under this condition, $\lim_{t\rightarrow\infty}x_{t}=\pi$ exists, $\pi>0$, and $\pi$ is independent of the initial probability distribution $x_{1}$. Moreover, $\lim_{k\rightarrow\infty}P^{k}=\pi e^{T}$, where, and in what follows, $e$ is the vector of all ones whose dimension can be determined by the context. We refer the reader to [15, 17] for the existing results on regular first order chains.
For the regular first order case, the limiting probability distribution is equivalent to the so-called stationary probability distribution. This appears to be the origin that the latter term has been used interchangeably in the literature on higher order chains. We, however, choose the former here because in the higher order case, the limiting probability distribution lacks an interpretation of stationarity as it comes to the evolution of the chain.
In light of the aforementioned results on first order chains, we now raise the question of whether parallel results can be established on higher order chains. As far as we can see, this question has been partially answered thus far by certain approximation schemes or a two-phase procedure under some stronger requirement, see a brief survey in Section 2. In this paper, we shall provide a more complete answer to this question. In particular, our results concern the exact limiting probability distribution without resorting to any approximation schemes. This is the main difference between our results and a vast majority of the relevant existing ones.
2 Background Material
Let us set the stage here first for the presentation of our results in the next section.
Let $m\geq 3$. We consider in the sequel an $(m-1)$th order Markov chain $X=\{X_{t}:t=1,2,\ldots\}$ on state space $S=\{1,2,\ldots,n\}$ with an $m$th order, $n$ dimensional transition tensor ${\cal P}=[p_{i_{1}i_{2}\ldots i_{m}}]$.
For $k\geq 1$, similar to (1.3), the $k$-step transition probability from states $(i_{2},\ldots,i_{m})$ to state $i_{1}$ is given by
@@ 2 Background Material
These $k$-step transition probabilities form a stochastic tensor too, which is denoted by ${\cal P}^{(k)}=[p^{(k)}_{i_{1}\ldots i_{m}}]$. The special case of $k=0$ leads to the identity tensor ${\cal I}=[\delta_{i_{1}i_{2}\ldots i_{m}}]$ — note that this identity tensor is different from those in [18, 23]. In addition, it is known that the $k$-step transition probabilities can be determined by the following recurrence formula [1, 9]:
$p^{(k+1)}_{i_{1}i_{2}\ldots i_{m}}=\sum_{j=1}^{n}p^{(k)}_{i_{1}ji_{2}\ldots i_{m-1}}p_{ji_{2}\ldots i_{m}},~{}k=0,1,2,\ldots.$ (2.1)
Using the symbol in [10], (2.1) can be written as a tensor product ${\cal P}^{(k+1)}={\cal P}^{(k)}\boxtimes{\cal P}$. Especially, the identity tensor satisfies ${\cal P}={\cal I}\boxtimes{\cal P}$. In this sense, ${\cal P}^{(k)}$ can be thought of as the $k$th power of $\cal P$. The operation $\boxtimes$ is clearly distributive, yet it is neither commutative nor associative.
Next, we recall the notions of irreducibility, ergodicity, and regularity of the transition tensor $\cal P$ (as well as the higher order chain $X$). These can be extended to nonnegative tensors too, but we shall restrict ourselves to the stochastic scenario here. Specifically, $\cal P$ is said to be :
@@
Obviously, regularity implies ergodicity. In the meantime, it is also known that ergodicity implies irreducibility but the reverse does not hold [9]. The latter is one of the differences between the higher order and the first order cases because irreducibility and ergodicity are known to be equivalent on a first order chain.
Now, we review a systematical way of converting the higher order chain $X$ to a first order one. Denote $T=\{i_{1}i_{2}\ldots i_{m-1}:i_{1},i_{2},\ldots,i_{m-1}\in S\}$. It is the collection of all multi-indices of length $m-1$. Such multi-indices are arranged via linear indexing [22]. When $m=4$ and $n=2$, for example,
@@
Clearly, the cardinality of $T$ is $N=n^{m-1}$. Consider random variables $Y_{t}=(X_{t},X_{t-1},\ldots,X_{t-m+2})$ that take values in $T$. Specifically, we define $Y_{t}=i_{1}i_{2}\ldots i_{m-1}$ whenever $X_{t}=i_{1},X_{t-1}=i_{2},\ldots,X_{t-m+2}=i_{m-1}$. Then, $Y=\{Y_{t}:t=m-1,m,\ldots\}$ is a first order chain on state space $T$ [15, 16]. It is also referred to as the reduced first order chain corresponding to $X$. The transition matrix $Q\in\mathbb{R}^{N\times N}$ of $Y$ is given by
@@
Observe that $q_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=p_{i_{1}i_{2}\ldots i_{m-1}j_{m}}$ whenever $i_{\ell}=j_{\ell}$ for all $\ell=2,3,\ldots,m-1$; otherwise, $q_{i_{1}i_{2}\ldots i_{m-1},j_{2}j_{3}\ldots j_{m}}=0$. Again, considering $m=4$ and $n=2$, we arrive at
@@
In $Q$, $p_{i_{1}i_{2}\ldots i_{m}}$ is on row $i_{1}+n(i_{2}-1)+\ldots+n^{m-2}(i_{m-1}-1)$ and column $i_{2}+n(i_{3}-1)+\ldots+n^{m-2}(i_{m}-1)$. As we shall see soon, there is an easier way of forming $Q$.
In the same manner as (1.3), the $k$-step transition probability of $Y$ from states $j_{2}j_{3}\ldots j_{m}$ to $i_{1}i_{2}\ldots i_{m-1}$ is given by
@@
It should be mentioned that problems regarding the higher order chain $X$ cannot always be fully addressed by resorting to the reduced first order chain $Y$. To explain this, we quote several examples here: (1) $Y$ may be non-ergodic even though $X$ is ergodic [10], (2) the mean first passage times of $X$ may not be obtained in a direct fashion from those of $Y$ [10], and (3) a recurrent state in $X$ may no longer be a part of any recurrent state in $Y$ [16]. The reduced first order chain, nevertheless, can still be a quite useful tool when dealing with the limiting probability distribution problem.
In the sequel, we denote [^2] In the first order case, this notation is consistent with that of the transition matrix. by $P\in\mathbb{R}^{n\times N}$ the mode-$1$ matricization [19] of the transition tensor $\cal P$. Simply put, $P$ consists of the frontal slices of $\cal P$ which are arranged side by side and ordered according to linear indexing. Taking the case $m=4$ and $n=2$, $P$ has the form
@@
We can also refer to the entries of $P$ in multi-index form as $p_{i_{1},i_{2}\ldots i_{m}}$. In $P$, $p_{i_{1}i_{2}\ldots i_{m}}$ is on row $i_{1}$ and column $i_{2}+n(i_{3}-1)+\ldots+n^{m-2}(i_{m}-1)$. In a similar vein, $P^{(k)}$ denotes the mode-$1$ matricization of ${\cal P}^{(k)}$. By convention, $P^{(0)}$ is the mode-$1$ matricization of the identity tensor $\cal I$, that is,
@@
Incidentally, the transition matrix $Q$ of the reduced first order chain is constructed in [26] using the product of a block diagonal matrix and a permutation matrix. By the mode-$1$ matricization of $\cal P$, the formulation of $Q$ can be simplified as
@@
More importantly, the mode-$1$ matricization of $\cal P$ can bridge the probability distribution of the higher order chain $X$ to that of the reduced first order chain $Y$. Specifically, let $y_{t}=x_{t,t-1,\ldots,t-m+2}\in\mathbb{R}^{N}$ be the probability distribution of $Y$ at time $t$, whose $i_{1}i_{2}\ldots i_{m-1}$th entry is given by
@@
for any $i_{1}i_{2}\ldots i_{m-1}\in T$. As usual, these entries are ordered by linear indexing. When $m=4$ and $n=2$, for example, we obtain
@@
By $x_{t}$ in (1.2) and $y_{t}$ above, it can be shown [21] that
@@
Formula (2.5) is exactly what the two-phase procedure in [26] is based on.
Now, with the foregoing background material at hand, we are ready to present a brief survey of the existing lines of inquiry regarding the limiting probability distribution of the higher order chain $X$.
The first (and the mainstream) approach assumes that for the reduced first order chain $Y$, if $\lim_{t\rightarrow\infty}y_{t}$ exists, then such a limit can be approximated by
$\underbrace{z\otimes_{K}\cdots\otimes_{K}z}_{m-1}$ (2.6)
for some $z\in\mathbb{R}^{n}$, where $\otimes_{K}$ stands for the Kronecker product. This approach is adopted in most existing works on higher order chains; see, for example, [1, 2, 3, 4, 5, 6, 8, 13, 14, 20, 21]. It leads to the interesting $Z$-eigenvector problem. It is known [21] that when $\cal P$ is irreducible, $z$ can be determined as some normalized positive $Z$-eigenvector. As pointed out in [26], however, the approximation in (2.6) may not capture the true nature of $\lim_{t\rightarrow\infty}y_{t}$.
The second approach [26] involves a two-phase power iteration scheme based on the connection between the higher order chain $X$ and its reduced first order chain $Y$. It assumes that $Q$ is regular, and so $\lambda=1$ is the only dominant eigenvalue of $Q$ on the unit circle and is simple. It follows that $\lim_{t\rightarrow\infty}y_{t}$ is the eigenvector, normalized using $\|\cdot\|_{1}$, of $Q$ associated with $\lambda$. Thus, the limiting probability distribution of $X$ can be determined by, for example, (2.5). As we shall demonstrate, however, the limiting probability distribution of $X$ may still exist despite $Q$ being non-regular.
The third approach [7] deals with the higher order chain $X$ by a certain approximation scheme to its transition probabilities. The development also hinges on a direct extension of the notion of recurrent states from the first order case to the higher order case. Such an extension has been shown [11] to be an issue. Incidentally, this issue manifests another difference between the first order and the higher order cases.
Motivated by the above works we have reviewed, especially by [26], our main goals here are to further investigate the properties of the actual (not approximated) limiting probability distribution of the higher order chain $X$ and to establish a new sufficient condition that guarantees the existence of such a limiting probability distribution.
3 Main Results
With the background material we have presented, let us begin this section with one example to illustrate that, in general, $\lim_{t\rightarrow\infty}x_{t}$ may depend on the initial probability distributions and, therefore, the limiting probability distribution may not exist.